HIGH RESOLUTION ELECTRONIC STARK SPECTROSCOPY OF P -AMINO BENZOIC ACID IN THE GAS PHASE^{a}

\underline{DIANE M. MITCHELL}, PHILIP J. MORGAN and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260.

Due to its biological relevance, hydrogen bonding ability, and “push-pull” structure, p−amino benzoic acid (PABA) has been the subject of several gas phase studies.^{b}^{c}^{d} Here, the structure of PABA in the S_{0} and S_{1} states has been determined using rotationally resolved electronic spectroscopy. Additionally, Stark spectra were taken while applying an electric field in order to determine the orientations and magnitudes of the permanent electric dipole moments of PABA in both electronic states. The results will be discussed.

aWork supported by NSF (CHE-0615755)

bG. Meijer, M. S. de Vries, H. E. Hunziker, and H. R. Wendt. J. Chem. Phys. 92, 7625 (1990).

cI. Compagnon, R. Antoine, D. Rayane, M. Broyer, and Ph. Dugourd. J. Phys. Chem. A. 107, 3036 (2003).

dY. He, C. Wu and W. Kong. J. Phys. Chem. A. 109, 2809 (2005).